# PLM-Based Discrete Diffusion Language Models with Entropy-Adaptive Gibbs Sampling

### v2

@@ E.1 LLM Evaluation
Table 10: RocStories LLM Evalution (LLM-c)
@@
E.2 Tables
We have appended the performance metrics of LLMs for reference. LLMs exhibit strong performance in tasks like QG, QQP, and Deontology, where datasets are more neat and contain fewer noises and abbreviations. Conversely, performance is lower on the Paradetox dataset, which includes a wide range of casual phrases and diverse perturbations. Provided that our model utilizes encoder-based PLMs, we observe that it demonstrates robustness to such perturbations compared to autoregressive models.
ParaDetox
PPL MAUVE SOME LLM-t VS(ngram) VS(emb) self-bleu distinct-1 distinct-2
GPT-2 389.147 0.503 0.717 0.02 3.925 2.640 0.429 0.312 0.748
GPT-3.5 104.375 0.175 0.888 0.074 3.098 1.915 0.652 0.390 0.835
GPT-4 78.979 0.125 0.879 0.18 3.214 1.906 0.592 0.412 0.841
DiffuSeq $\geq 1k$ 0.683 0.703 0.03 2.059 1.465 0.841 0.410 0.820
Diffusion-Bert 775.928 0.737 0.716 0.09 3.101 2.058 0.599 0.424 0.826
DINOISER 124.797 0.255 0.767 0.1 2.287 2.174 0.981 0.211 0.486
SEDD-small $\geq 1k$ NA 0.664 NA 4.746 4.063 0.119 0.451 0.846
Diffusion-EAGS 69.5 0.773 0.796 0.01 4.755 3.659 0.126 0.475 0.834
Table 11: ParaDetox Dataset Generation : We use LLM-evaluation approach for measuring toxicity, denoted as LLM-t
Deontology
PPL MAUVE SOME VS(ngram) VS(emb) self-bleu distinct-1 distinct-2
GPT-2 91.962 0.131 0.860 3.665 3.126 0.425 0.474 0.874
GPT-3.5 52.401 0.393 0.904 4.632 3.650 0.186 0.434 0.855
GPT-4 72.329 0.465 0.921 4.530 3.286 0.191 0.425 0.865
Table 12: Deontology Dataset Generation
QQP
PPL MAUVE SOME VS(ngram) VS(emb) self-bleu distinct-1 distinct-2
GPT-2 66.270 0.112 0.754 3.886 2.566 0.423 0.344 0.787
GPT-3.5 55.275 0.708 0.874 2.781 1.603 0.691 0.365 0.814
GPT-4 66.121 0.814 0.877 2.981 1.651 0.673 0.423 0.866
Table 13: QQP Dataset Generation
QG
PPL SOME MAUVE VS(ngram) VS(emb) self-bleu distinct-1 distinct-2
GPT-2 124.831 0.759 0.175 4.564 3.130 0.176 0.210 0.629
GPT-3.5 57.539 0.808 0.0184 3.174 2.090 0.612 0.282 0.739
GPT-4 75.435 0.822 0.0171 3.819 2.069 0.444 0.311 0.773
Table 14: QG Dataset Generation
Appendix F Well-Generated Output Examples
F.1 ParaDetox
Table 15: ParaDetox generation examples
@@ F.1 ParaDetox
Table 16: As the dataset is designed to eliminate problematic words while preserving semantics, our model adheres to these guidelines.
F.2 Deontology
Table 17: Deontology generation examples
@@ F.2 Deontology
F.3 QQP
Table 18: QQP generation examples
@@ F.3 QQP
F.4 QG
Table 19: QG generation examples
@@
F.5 RocStories
Table 20: RocStories generation examples
@@ F.5 RocStories
Appendix G Details on Text Augmentation Using GPT models
G.1 GPT-3.5turbo ~ GPT-4-Omni
@@ Appendix G Details on Text Augmentation Using GPT models
Appendix H Details on CDLMs
H.1 Experimental Details
@@ H.2 Results Interpretations
Appendix I Details on Translation Results
I.1 Comparison Between Easy-First and Our Proposed Method
@@ I.2 Experimental Details
Diffuseq-v2 For Diffuseq-v2, we employ the vocabs of mBERT and choose 128 as max length for ende translation. Other settings are identical as in Appendix H.1.
@@
I.3 Experimental Results
I.3.1 NAR Transformer, DisCo
@@ I.3.2 Diffuseq-v2
Appendix J Entropy Flow
In Figure 4, we illustrate the tendency of the sequential sum of entropy for various discrete generation processes. The changes of entropy during the generation process in Diffusion-EAGS, represented by the yellow line, show that our model effectively follows a gradual decrease in entropy, mirroring the inverse trend of the training process. This gradual change in entropy facilitates successful DDLM training, which results in superior text quality performance compared to other diffusion models, as demonstrated in Tables 3, 2, and 4.

### v3

@@ H.1 LLM Evaluation
Table 16: ParaDetox LLM Evalution (LLM-t)
@@
Table 17: RocStories LLM Evalution (LLM-c)
Appendix I Well-Generated Output Examples
Generated examples of Paradetox are provided in Table 18, Deontology in Table 19, QQP in Table 20, QG in Table 21, and RocStories in Table 22.
@@ Appendix I Well-Generated Output Examples
Table 18: ParaDetox generation examples
@@
Table 19: Deontology generation examples
@@
Table 20: QQP generation examples
@@
Table 21: QG generation examples
@@
Table 22: RocStories generation examples
Appendix J Details on Text Augmentation Using GPT models
J.1 GPT-3.5turbo ~ GPT-4-Omni
@@
Appendix K Details on CDLMs
K.1 Experimental Details
@@
Appendix L Details on Translation Results
L.1 Datasets & Observations
[…]
Interestingly, the output of the pre-trained NLLB model (NLLB-naive-600M, not finetuned) reveal that neural network-based metrics are susceptible to the interference problem, specifically translated by other languages, even though we provide the language specific token. While such issues result in lower BLEU scores, COMET and XCOMET often interpret them as semantically coherent, indicating a potential direction for future work to improve translation evaluation metrics. Despite these phenomena, a performance gap between translation models and DDLM remains. This suggests that future research should address the semantic capabilities of diffusion models to help bridge this gap.
L.2 Comparison Between Easy-First and Our Proposed Method
@@
Diffuseq-v2 For Diffuseq-v2, we employ the vocabs of mBERT and choose 128 as max length for ende translation. Other settings are identical as in Appendix K.1.
@@
L.4 Experimental Results
L.4.1 NAR Transformer, DisCo